OpenStudy (anonymous):

medal and fan What is the y-intercept and asymptote of this function? f(x) = 5^x A.(0,1), y=0 B.(1,0), y=0 C.(0,1), x=0 D.(1,0), x=0

OpenStudy (anonymous):

C.(0,1), x=0

OpenStudy (anonymous):

wrong its a

OpenStudy (anonymous):

The y-intercept is when x = 0 f(0) = 5^0 = 1 (0,1) The function is asymptotic to y = 0. f(-1) = 1/5 f(-2) = 1/25 f(-3) = 1/125 As x gets smaller and smaller, the function gets closer and closer to 0, but does not cross it.

OpenStudy (anonymous):

right